# Rotational Properties of Inverted Hybrid Stars

### v1

@@ V Back-Bending
A typical manifestation of the possibility of a phase transition in the core of neutron stars is the phenomena known as “back-bending” [44]. This phenomenon can be comprehended by considering the star’s contractions during its rotational evolution, which reduces the moment of inertia, in conjunction with the softening of the EOS that accompanies the phase transition. These combined factors lead to an increase in the star’s rotational frequency, despite a reduction in its total angular momentum during spin-down. This behavior can be distinctly identified by introducing an additional axis, representing the rotational frequency, to Fig. 6. The resulting graphic, depicted in Fig. 8, clearly demonstrates an increase in rotational frequency corresponding to an increase in stellar density during the star’s rotational evolution.
Figure 8: 3D representation of the stellar sequences shown in Fig. 6, with the z-axis representing the rotational frequency.
@@
Figure 9: Angular momentum as a function of frequency for constant baryonic mass sequences of EOS B.
@@
The results depicted in Fig. 9 exhibit intriguing behavior. While they retain the expected properties associated with the previously studied back-bending phenomenon (an increase in rotational frequency corresponding to a decrease in angular momentum), the shape of the $J$-$\Omega$ curve is markedly different. Traditional back-bending curves typically display an ‘S’ shape, hence the name ‘back-bending’. However, for the cross stars studied here, we observe a much sharper bend in the curve, more akin to a ‘back-kink’. We attribute this behavior to the nature of the inverted hybrid stars under study and the abrupt density change in nature when the low-pressure $ud$QM transitions to hadronic matter at higher pressures. Although we currently have limited observational data regarding the angular momentum of compact stars, future observations of this nature could potentially provide empirical evidence for the model studied here.
VI Conclusions
@@
In this research, we built upon the work of [17, 18, 20] which has put forth the intriguing possibility that quark matter can be more stable than hadronic matter at low pressure and be less stable at higher pressure. This provocative possibility gives rise to inverted hybrid stars or cross stars (CrSs) by other names, in which their outer layer is composed of quark matter while the higher-density core is made of hadrons, wherein the quarks are confined.
@@
The studies presented in [17, 18] comprehensively examine the characteristics of cross stars, such as mass, radius, tidal deformability, and oscillations. Our research aims to extend their findings by investigating the attributes of rotating CrSs, including the potential evolutionary trajectories of objects with constant baryonic mass. We find that such studies are garnering renewed interest, especially considering the possibility of post-merger objects with high angular momentum. We have then analyzed three benchmark models for CrSs, each representing a different “depth” of the $ud$QM to hadronic matter phase transition: deep (EOS A), intermediate (EOS B), and shallow (EOS C). All equations of state presuppose the APR EOS for the hadronic matter. We have computed various families of rotating neutron stars across all models to gain a comprehensive understanding of their properties.
@@
Our research reveals that CrSs rotating at their Kepler frequencies typically exhibit a higher mass and a larger circumferential radius, as anticipated. We have also observed an interesting behavior regarding twin configurations, where twin stars are compact objects with identical mass but differing radii, a characteristic often observed in hybrid stars. Our results indicate that in configurations with high angular momentum, such as those in the Kepler sequence, the likelihood of potential twin configurations increases significantly. Moreover, EOS B and EOS C start to exhibit twin stars within their Kepler sequences, despite not presenting such configurations in the non-rotating case.
@@
We have also devoted considerable effort to studying sequences of constant baryonic mass. These sequences are characterized by different stellar configurations that share the same baryonic mass but exhibit varying angular momentum. These configurations reach their extremes at the Kepler sequence, beyond which no hydrostatically stable configurations exist, and at the spherical sequence, which represents the minimum possible angular momentum. Sequences of constant baryonic mass between these two extremes represent potential paths of rotational evolution that stars may follow as they transition from high angular momentum to spherical, non-rotating configurations. We have shown that although we find a sequence of constant baryon mass connecting stars at the Kepler frequency to their non-rotating counterparts, not all stars in these sequences are viable. While all of them satisfy the hydrostatic equilibrium conditions, some of them are not stable against oscillations. These unstable configurations occur due to the onset of phase transition during the spin-down, which happens as the central density increases during the spin-down, triggering the phase transition within the star. This leads to the possibility of a mini-collapse, occurring in the spin-down of stars with high angular momentum, such as those studied in Ref. [44]. We have estimated that for the current model, these mini-collapses are associated with the release of $\sim 10^{50}$ erg of energy, part of which would be released in the form of electromagnetic and/or gravitational radiation.
@@
We have further investigated the properties of constant baryonic mass during spin-down sequences, particularly the phenomena of “back-bending”. This phenomenon is characterized by an increase in rotational frequency that coincides with a reduction in angular momentum for stars nearing the onset of phase transition. Typically observed in stars undergoing phase transition during spin-down, this phenomenon arises from the need for a rapid increase in rotational frequency to compensate for a substantial decrease in the moment of inertia during the phase transition. This is usually manifested as an ‘S’ shape in the angular momentum graph. However, due to the unique nature of CrSs, which have a quark phase in the outer layers, the back-bending for these stars is exhibited in a rather distinctive way, behaving more akin to what can be described as a ‘back-kink’ due to its sharp shape. Although admittedly remote given our limited observational data on the angular momentum and moment of inertia of neutron stars, there may be a possibility of distinguishing between CrSs and ordinary hybrid stars if more observational data on the relevant quantities become available in the future.
@@
With this work, our objective has been to build upon previous studies of inverted hybrid stars by incorporating the significant aspect of rotation. Our research has unveiled intriguing phenomena that occur in this new type of compact star when angular momentum is taken into account, some of which could potentially serve as observational markers with the emergence of new data. We believe that our work enriches the studies presented by [17, 20], introducing an additional and captivating layer of complexity. We intend to delve deeper into the fascinating concept of CrSs by investigating differential rotation and magnetic fields, a study that is currently in progress and will be detailed in our future works.

### v2

@@ V Back-Bending
A typical manifestation of the possibility of a phase transition in the core of neutron stars is the phenomena known as “back-bending” Haensel et al. (2016). This phenomenon can be comprehended by considering the star’s contractions during its rotational evolution, which reduces the moment of inertia, in conjunction with the softening of the EOS that accompanies the phase transition. These combined factors lead to an increase in the star’s rotational frequency, despite a reduction in its total angular momentum during spin-down. This behavior can be distinctly identified by introducing an additional axis, representing the rotational frequency, to Fig. 6. The resulting graphic, depicted in Fig. 8, clearly demonstrates an increase in rotational frequency corresponding to an increase in stellar density during the star’s rotational evolution.
Figure 8: 3D representation of the stellar sequences shown in Fig. 6, with the $z$-axis representing the rotational frequency. Solid black and red lines represent spherical and Kepler sequences, respectively. Cyan, purple and green curves represent the constant baryonic mass sequences. Also shown in the graph are the projection on the three perpendicular planes, as well as droplines connecting the 3D curves to the $x-y$ plane.
@@
Figure 9: Angular momentum as a function of frequency for constant baryonic mass sequences of EOS B. Each colored curve follow constant baryonic mass as indicated in the figure. Orange shaded region represents unstable configurations.
@@
The results depicted in Fig. 9 exhibit intriguing behavior. Although they retain the expected properties associated with the previously studied back-bending phenomenon (an increase in rotational frequency corresponding to a decrease in angular momentum), the shape of the $J$-$\Omega$ curve is markedly different. Traditional back-bending curves typically show an ‘S’ shape, hence the name ‘back-bending’. However, for the inverted hybrid stars studied here, we observe a much sharper bend in the curve, more akin to a ‘back-kink’. We attribute this behavior to the nature of the inverted hybrid stars under study and the abrupt density change in nature when low-pressure $ud$ QM transitions to hadronic matter at higher pressures. We note that previous studies have also identified a similar behavior as the ‘back-kink,’ as observed in Zdunik and Gourgoulhon (2006); Bhattacharyya et al. (2005). In these works, ordinary hybrid stars also exhibit a similar ‘back-kink,’ which can be attributed to the significant discontinuity in density at the phase transition. A key advantage of the inverted hybrid star model is that, even with a sharp first-order phase transition and large density discontinuities, its $M_{\rm TOV}$ is more likely to reach the desired $2M_{\odot}$ target (see Table I of Ref. Zhang et al. (2024)). This allows for the existence of a more massive hybrid object ($M>2M_{\odot}$) that exhibits a more pronounced back-kink behavior than normal hybrid stars. Although we currently have limited observational data regarding the angular momentum of compact stars, future observations of this nature could potentially provide empirical evidence for the model studied here.
VI Braking Index
A relevant observable quantity for pulsars is the braking index, which is defined in terms of the observable frequency ($\Omega$) and its first and second derivatives ($\dot{\Omega},\ddot{\Omega}$) and is written as Manchester and Taylor (1977); Lyne and Graham-Smith (2012); Blandford and Romani (1988); Melatos (1997)
$n=\frac{\Omega\ddot{\Omega}}{\dot{\Omega}^{2}}.$ (12)
In the slow rotation approximation, the spin-down of a neutron star is normally assumed to take the form of a power law. In the ideal case where the only braking is due to the magnetic field, one obtains $n=3$. On the other hand, if the braking is due to gravitational wave emission one gets $n=5$ Lyne and Graham-Smith (2012).
A more accurate description for the braking index of a relativistic, rapidly rotating neutron star is found in Spyrou and Stergioulas (2002) which can be written, for the case in which the spin-down is solely due to the magnetic field as
$n(\Omega)=3-\frac{2I^{\prime}\Omega+I^{\prime\prime}\Omega^{2}}{I+I^{\prime}\Omega},$ (13)
where a prime indicates a derivative with respect to the frequency ($\Omega$).
The braking index can be considered an interesting way of probing the interior of pulsars, as observations can be compared with predicted values to provide hints about the stellar interior. We begin our analysis by studying the moment of inertia of the three representative stars studied in the previous section ($M_{B}=1.80,2.05$ and $2.31$ M). The dependence of the moment of inertia as a function of frequency is shown in Fig. (10).
Figure 10: Moment of inertia as a function of rotation frequency for three representative stars of EOS B. Each curve represents the indicated baryonic mass sequence, as indicated on the figure. The orange-shaded region represents unstable configurations.
We note that all three stars exhibit similar qualitative behavior, with the moment of inertia smoothly growing for the stars with lower frequencies (higher densities) which have both hadronic and quark phases. The moment of inertia thus turns back at a maximum frequency, only to turn back again, in an effect mimicking the back-bending studied in the previous section (which in reality is a result of this behavior for the moment of inertia in stars with phase transition).
Such behavior indicates that we can expect wildly different values for the braking index of stars with this equation of state, much like those stars studied in Spyrou and Stergioulas (2002). To further investigate the braking index of these stars we now focus only on the $M_{B}=1.80$ M, as the same qualitative behavior is found for the other stars. We show in Fig. (11) the star’s moment of inertia, with 4 separate regions that will be analyzed.
Figure 11: Moment of inertia as a function of rotation frequency the star with $M_{B}=1.80$ M of EOS B. Also shown are four different regions which will be used to study the star’s braking index.
Using the moment of inertia as a function of rotational frequency, shown in Fig. (11), we use eq. (13) to calculate the braking index for the star in the four indicated regions. This result is shown in Fig. 12
Figure 12: Braking index for the $M_{B}=1.80$ M divided in the the four regions indicated in Fig.11. The insets indicate the corresponding moment of inertia section.
Fig. 12 indicates that for lower frequencies, thus higher densities and lower deformations, the star has a breaking index near the canonically expected value of 3. As the frequencies increase and deformations start to take place, the non-linear aspects of the moment of inertia evolution become noticeable and we have the braking index taking large negative or positive values. These results are somewhat typical of stars with first-order phase transition in their interior, as discussed in Spyrou and Stergioulas (2002), and may be useful for comparison with observed pulsars that may exhibit unusual values for the braking index, as this may be an indication of a phase transition to exotic phases.
VII Universal Relations
There has been great interest in studying EOS insensitive relations for rotating neutron stars, the so-called universal relations. Many papers have devoted great efforts to finding such relations Breu and Rezzolla (2016); Bozzola et al. (2019); Khadkikar et al. (2021). We now discuss a few of these relations, particularly those associated with supramassive configurations, as well as the relationship between the moment of inertia and quadrupole moment ($I-Q$), which holds particular relevance for the study of merger events.
We now discuss a universal relation for the supramassive configurations, which are stars more massive than their non-rotating counterparts. In order to do that, we must analyze several sequences of constant angular momentum ($J$) for the three EOS’s studied. Such sequences are shown in Figs. 13, 14 and 15.
Figure 13: Sequence of stars with constant angular momentum for EOS A. The solid black line represents spherically symmetric stars with zero angular momentum, whereas the red curve represents the Kepler sequence, in which stars rotate at their maximum possible frequency. Dashed lines represent sequences with constant angular for a range of values between zero and maximum allowed rotation.
Figure 14: Same as Fig. 13 except for EOS B.
Figure 15: Same as Fig. 13 except for EOS C.
We now turn our attention to the maximum mass configuration for the sequences depicted in Figs. 1315, which we define as $M_{\uparrow}$. For the spherically symmetric sequence, this is nothing less than the maximum TOV mass ($M_{max}^{TOV}$), whereas for the Kepler sequence this star represents the supramassive limit, i.e. the upper limit mass that a (rigidly) rotating star may have. The authors of Breu and Rezzolla (2016); Bozzola et al. (2019) have found, for hadronic stars, a universal relation between $M_{\uparrow}$ normalized by $M_{max}^{TOV}$ and the corresponding star’s angular momentum $J$, normalized by the maximum possible angular momentum ($J_{Max}$), which is written as
$\frac{M_{\uparrow}}{M_{Max}^{TOV}}=1+0.1316\left(\frac{J}{J_{Max}}\right)^{2}+0.0711\left(\frac{J}{J_{Max}}\right)^{4}$ (14)
We perform a similar study, now considering the models under investigation in this work. In order to do that, we have numerically found the maximum mass configuration for the different constant angular momentum sequences (normalized by $J_{Max}$) , normalizing it by their corresponding $M_{max}^{TOV}$. With this quantity, we have found that the best fit curve for inverted hybrid stars is given by
$\frac{M_{\uparrow}}{M_{Max}^{TOV}}=1+a\left(\frac{J}{J_{Max}}\right)^{2}-b\left(\frac{J}{J_{Max}}\right)^{4},$ (15)
where the constants assume the values $a=0.3186\pm 0.017$ and $b=0.0886\pm 0.024$.
Our results are shown graphically in Fig. 16, where besides the numerical data, we also show the best-fit curve as well as the confidence and prediction band of $95\%$.
Figure 16: Normalized gravitational mass of the top mass as a function of the corresponding normalized stellar angular momentum, for the three different EOSs studied. Also shown are the best-fit curve and the 95% confidence and prediction bands. For comparison purposes, we also show the curve for Eq. (14).
The results depicted in Fig. 16 indicate that most likely IHSs also satisfy universal relations, at the very least for the supramassive limit, although this would need to be confirmed by a more exhaustive study with a larger set of EOS for IHSs and also investigating other universal relations.
Now we must also spend some time discussing the relationship between the star’s moment of inertia ($\hat{I}$) and the spin-induced quadrupole moment ($\hat{Q}$). The authors of Yagi and Yunes (2013a, b) have shown that, for slowly rotating neutron stars, there is a universal relation connecting $\hat{I}=I/M^{3}$ and $\hat{Q}=-Q/M^{3}a^{2}$ (where $a=J/M^{2}$) given by
$\begin{split}\ln{\hat{I}_{YY}}=1.35+0.697\ln{\hat{Q}}-0.143\ln{\hat{Q}^{2}}+\\ 9.94\times 10^{-2}\ln{\hat{Q}^{3}}-1.24\times 10^{-2}\ln{\hat{Q}^{4}}.\end{split}$ (16)
The work presented in Doneva et al. (2013) has extended this UR study to rapidly rotating neutron stars, and found a breakdown of universality in the rapidly rotating case. The authors of Chakrabarti et al. (2014) made further refinements showing that a much better universality is achieved if certain dimensionless parameters characterizing the magnitude of rotation are used, such as $\tilde{f}=20R\Omega$, where $R$ is the equatorial radius. For this part of our study we follow the approach of Chakrabarti et al. (2014) and calculate the relationship between $\hat{I}$ and $\hat{Q}$ for constant sequences $\tilde{f}$. This result is shown in Fig. 17, where we also compare the $\hat{I}-\hat{Q}$ relation with the fit obtained by Yagi and Yunes (2013a, b).
Figure 17: $\hat{I}$ as a function of $\hat{Q}$ for stars with different values of the parameter $\tilde{f}$ as indicated by the figure. Different symbols denote different EOS, with squares representing EOS A, circles EOS B, and diamonds EOS C. Also shown is the relation found by Yagi and Yunes (2013a, b) for slowly rotating neutron stars
We can surmise by the results of Fig. 17 that IHSs behave similarly to the other models studied in reference Doneva et al. (2013), where it was shown that rapidly rotating stars deviate from the fit found in Yagi and Yunes (2013a, b), especially for higher values of $\hat{Q}$.
Ref. Chakrabarti et al. (2014) extended the $\hat{I}-\hat{Q}$ fitting relation eq. (16) with an expansion of the parameter $\tilde{f}$ to achieve better universality in the rapidly rotating case. Therefore, we now compare the numerical $\hat{I}-\hat{Q}$ found for IHSs (Fig. 17) with the fit proposed in Chakrabarti et al. (2014) (labeled here $I_{CH}$), by showing the relative deviation between these quantities in Fig. 18.
Figure 18: Relative deviation of the calculated $\hat{I}$ to the fit $\hat{I}_{CH}$ for stars with different values of the parameter $\tilde{f}$ as indicated by the figure. Different symbols denote different EOS, with squares representing EOS A, circles EOS B, and diamonds EOS C.
The results shown in Fig. 18 suggest that IHSs adhere to the universal behavior outlined in Chakrabarti et al. (2014), particularly for smaller values of $\hat{Q}$ and stars with low $\tilde{f}$. However, there are relatively larger discrepancies for higher values of $\hat{Q}$. These findings diverge from those reported by Paschalidis et al. (2018), where the $\hat{I}-\hat{Q}$ relation for different (ordinary) hybrid stars revealed a smaller relative discrepancy.
We speculate that this variation arises from the absence of a nuclear crust in IHSs, which have an outer region composed of quark matter, unlike ordinary neutron and hybrid stars. This lack of a crust may explain the observed discrepancies in high $\hat{Q}$ stars, which are associated with smaller central densities where low-density quark matter plays a more significant role.
To shed further light on this behavior, we plan to investigate the properties of the $\hat{I}-\hat{Q}$ relation for bare quark stars compared to neutron stars and IHSs in future studies.
VIII Conclusions
@@
In this research, we built upon the work of Zhang and Ren (2023); Zhang et al. (2024); Holdom et al. (2018) which has put forth the intriguing possibility that quark matter can be more stable than hadronic matter at low pressure and be less stable at higher pressure. This provocative possibility gives rise to inverted hybrid stars or cross stars by other names, in which their outer layer is composed of quark matter while the higher-density core is made of hadrons, wherein the quarks are confined.
@@
The studies presented in Zhang and Ren (2023); Zhang et al. (2024) comprehensively examine the characteristics of cross stars, such as mass, radius, tidal deformability, and oscillations. Our research aims to extend their findings by investigating the attributes of rotating IHSs, including the potential evolutionary trajectories of objects with constant baryonic mass. We find that such studies are garnering renewed interest, especially considering the possibility of post-merger objects with high angular momentum. We have then analyzed three benchmark models for IHSs, each representing a different “depth” of the $ud$QM to hadronic matter phase transition: deep (EOS A), intermediate (EOS B), and shallow (EOS C). All equations of state presuppose the APR EOS for the hadronic matter. We have computed various families of rotating neutron stars across all models to gain a comprehensive understanding of their properties.
@@
Our research reveals that IHSs rotating at their Kepler frequencies typically exhibit a higher mass and a larger circumferential radius, as anticipated. We have also observed an interesting behavior regarding twin configurations, where twin stars are compact objects with identical mass but differing radii, a characteristic often observed in hybrid stars. Our results indicate that in configurations with high angular momentum, such as those in the Kepler sequence, the likelihood of potential twin configurations increases significantly. Moreover, EOS B and EOS C start to exhibit twin stars within their Kepler sequences, despite not presenting such configurations in the non-rotating case.
@@
We have also devoted considerable effort to studying sequences of constant baryonic mass. These sequences are characterized by different stellar configurations that share the same baryonic mass but exhibit varying angular momentum. These configurations reach their extremes at the Kepler sequence, beyond which no hydrostatically stable configurations exist, and at the spherical sequence, which represents the minimum possible angular momentum. Sequences of constant baryonic mass between these two extremes represent potential paths of rotational evolution that stars may follow as they transition from high angular momentum to spherical, non-rotating configurations. We have shown that although we find a sequence of constant baryon mass connecting stars at the Kepler frequency to their non-rotating counterparts, not all stars in these sequences are viable. While all of them satisfy the hydrostatic equilibrium conditions, some of them are not stable against oscillations. These unstable configurations occur due to the onset of phase transition during the spin-down, which happens as the central density increases during the spin-down, triggering the phase transition within the star. This leads to the possibility of a mini-collapse, occurring in the spin-down of stars with high angular momentum, such as those studied in Ref. Haensel et al. (2016). We have estimated that for the current model, these mini-collapses are associated with the release of $\sim 10^{51}$ erg of energy, part of which would be released in the form of electromagnetic and/or gravitational radiation.
@@
We have further investigated the properties of constant baryonic mass during spin-down sequences, particularly the phenomena of “back-bending”. This phenomenon is characterized by an increase in rotational frequency that coincides with a reduction in angular momentum for stars nearing the onset of phase transition. Typically observed in stars undergoing phase transition during spin-down, this phenomenon arises from the need for a rapid increase in rotational frequency to compensate for a substantial decrease in the moment of inertia during the phase transition. This is usually manifested as an ‘S’ shape in the angular momentum graph. However, due to the unique nature of IHSs, which have a quark phase in the outer layers, the back-bending for these stars is exhibited in a rather distinctive way, behaving more akin to what can be described as a ‘back-kink’ due to its sharp shape. Although admittedly remote given our limited observational data on the angular momentum and moment of inertia of neutron stars, there may be a possibility of distinguishing between IHSs and ordinary hybrid stars if more observational data on the relevant quantities become available in the future.
Furthermore, we have calculated the braking index of IHSs during spin-down. Our results indicate that for low-frequency stars, when their shape is nearly spherical, the braking index approximates the canonical value of three, as expected. For higher frequencies, where the star’s shape and moment of inertia deviate significantly from those of spherical stars, the effects of back bending become apparent on the braking index. This index substantially increases (in absolute value) in the back-bending region. These observations could potentially serve as evidence for probing the star’s internal composition.
We have also explored whether universal relations for the supramassive mass can be found across all studied models. This relation is significant as they relate the maximum mass of a rigidly rotating star to the TOV mass limit and the star’s angular momentum. We have identified such a connection with a relatively high level of numerical confidence. We believe that by investigating more IHSs EOS models, we can achieve an even more precise correlation. Our future plans include continuing this investigation with more microscopic models for IHSs and expanding our search for universal relations for IHSs.
Furthermore, concerning universal relations, our studies on the $\hat{I}-\hat{Q}$ relation of IHSs have demonstrated that they exhibit similar behaviors to the universal relations found for rapidly rotating stars, albeit only for low values of $\hat{Q}$. Beyond these values, discrepancies emerge. We attribute this behavior to the quark matter crust of IHSs and plan to investigate this phenomenon further in future research.
@@
With this work, our objective has been to build upon previous studies of inverted hybrid stars by incorporating the significant aspect of rotation. Our research has unveiled intriguing phenomena that occur in this new type of compact star when angular momentum is taken into account, some of which could potentially serve as observational markers with the emergence of new data. We believe that our work enriches the studies presented by Zhang and Ren (2023); Holdom et al. (2018), introducing an additional and captivating layer of complexity. We intend to delve deeper into the fascinating concept of IHSs by investigating differential rotation and magnetic fields, a study that is currently in progress and will be detailed in our future works.